# Amortized Variational Transdimensional Inference

### v1

@@ F.4 Simulation design
We generate 10 i.i.d. complete data sets of length $n_{\operatorname{max}}=2^{10}$ from the above process. The experiment compares data size against the metrics from Section˜F.2. The range of data sizes are $n=16,32,4,128,256,512,1024$, where $n<n_{\operatorname{max}}$ simply takes the first $n$ samples.
@@
F.5 Real data example
For VTI, we chose a penalized structural model prior $p(m)$ where $\lambda=200$. We set the number of hidden nodes per edge to $H=5$ and include the bias terms, i.e. $\beta=1$.
For DAGMA non-linear, we use 10 hidden nodes per edge and no bias term.
F.6 DAG Model indicator construction: Lehmer Code Decoding
@@
The permutation $\varpi$ is stored as a one-hot $\bm{P}\in\{0,1\}^{N_{d}\times N_{d}}$ with $\bm{P}_{r,i}=1$ iff row $r$ is chosen at column $i$.
@@
2:  $\bm{P}\leftarrow\mathbf{0}_{\,\text{bs}\times N_{d}\times N_{d}}$
@@
6:  $\text{Used}\leftarrow\displaystyle\sum_{c=1}^{i-1}\bm{P}[:,:,c]$
@@
9:  $\bm{P}\bigl[\text{Idx}_{[:,0]},\,\text{Idx}_{[:,1]},\,i\bigr]\leftarrow\text{reshape}\!\bigl(\text{OneHot},\,{-}1\bigr)$
@@
11:  $\text{Used}\leftarrow\displaystyle\sum_{c=1}^{N_{d}-1}\bm{P}[:,:,c]$
@@
13:  $\bm{P}\bigl[\text{Last}_{[:,0]},\,\text{Last}_{[:,1]},\,N_{d}\bigr]\leftarrow 1$
14:  return $\bm{P}$
F.7 Model identifier for directed acyclic graphs
We encode a permutation matrix $\bm{P}\in\{0,1\}^{N_{d}\times N_{d}}$ using a compressed Lehmer code consisting of $N_{d}-1$ categorical variables $\{\rho^{\operatorname{cat}}_{1},\ldots,\rho^{\operatorname{cat}}_{N_{d}-1}\}$. Here $\rho^{\operatorname{cat}}_{i}$ has $N_{d}-i+1$ outcomes.
@@ F.7 Model identifier for directed acyclic graphs
Given $\bm{P}$ we form an upper-triangular mask $\bm{U}\in\{0,1\}^{N_{d}\times N_{d}}$ with zero diagonal. Each entry above the diagonal ($i<j$) is a Bernoulli variable, so $\bm{U}$ flattens to $\tfrac{N_{d}(N_{d}-1)}{2}$ bits. The adjacency matrix is $\bm{A}=\bm{P}^{\top}\bm{U}\bm{P}$, giving a DAG.
@@
The identifier $\{\rho^{\operatorname{cat}}_{1},\ldots,\rho^{\operatorname{cat}}_{N_{d}-1},\bm{U}_{\mathsf{binary}}\}$ is modelled autoregressively by a single $\text{MADE}^{+}$ network, yielding $\bm{A}=\bm{P}^{\top}\bm{U}\bm{P}$ upon sampling.
@@
$p(\bm{P},\bm{U}\mid\gamma)$ $=\frac{1}{N_{d}!}\;\frac{1}{2^{\frac{N_{d}(N_{d}-1)}{2}}}\;\exp\!\bigl(-\gamma\,\mathrm{nEdges}(\bm{U})\bigr),$ (59)
$\mathrm{nEdges}(\bm{U})$ $=\sum_{i<j}\bm{U}_{ij},$ (60)
$\log p$ $=-\log(N_{d}!)-\frac{N_{d}(N_{d}-1)}{2}\log 2-\gamma\,\mathrm{nEdges}(\bm{U}).$ (61)
@@
F.8 Neural probability mass function for model indicators over large spaces: $\text{MADE}^{+}$
To represent a distribution over binary strings, we use the Masked Autoencoder for Density Estimation (MADE) (Germain et al., , 2015) implementation found in the Durkan et al., (2020) repository. To represent a more complex discrete distribution such as that required by the $\bm{P},\bm{U}$ representation of a directed acyclic graph, we apply a simple extension to this architecture to allow us to vary the output dimension multiplier. For presentational clarity we call this extension $\text{MADE}^{+}$. The key change in $\text{MADE}^{+}$is the introduction of a per-dimension output multiplier function r(i)\,r(i)\, that determines how many parameters are emitted for the $i$-th input dimension in the autoregressive factorization.

### v2

@@ G.4 Simulation design
We generate 10 i.i.d. complete data sets of length $n_{\operatorname{max}}=2^{10}$ from the above process. The experiment compares data size against the metrics from Appendix G.2. The range of data sizes are $n=16,32,4,128,256,512,1024$, where $n<n_{\operatorname{max}}$ simply takes the first $n$ samples.
@@
In the DAGMA setup, a sweep of the regularization tuning parameter $\lambda$ was conducted for each dataset. The resulting adjacency matrix with the closest number of active edges to the data-generating graph was selected. This resulted in a higher-than-usual score for DAGMA results in the simulation study when compared to other methods. For DiBS/DiBS+, the inference ran for $5,000$ steps over $10$ “particles” (each an individual Stein variational gradient descent optimization, see Lorch et al. [2021]). JSP-GFN was configured to use a batch size of $1024$ over $50,000$ iterations.
G.5 Real data example
[…]
For DAGMA non-linear, DiBS/DiBS+, and JSP-GFN, we use 10 hidden nodes per edge and no bias term.
G.6 DAG Model indicator construction: Lehmer Code Decoding
@@
The permutation $\varpi$ is stored as a one-hot $\mathbf{P}\in\{0,1\}^{N_{d}\times N_{d}}$ with $\mathbf{P}_{r,i}=1$ iff row $r$ is chosen at column $i$.
@@
2:  $\mathbf{P}\leftarrow\mathbf{0}_{\,\text{bs}\times N_{d}\times N_{d}}$
@@
6:  $\text{Used}\leftarrow\displaystyle\sum_{c=1}^{i-1}\mathbf{P}[:,:,c]$
@@
9:  $\mathbf{P}\bigl[\text{Idx}_{[:,0]},\,\text{Idx}_{[:,1]},\,i\bigr]\leftarrow\text{reshape}\!\bigl(\text{OneHot},\,{-}1\bigr)$
@@
11:  $\text{Used}\leftarrow\displaystyle\sum_{c=1}^{N_{d}-1}\mathbf{P}[:,:,c]$
@@
13:  $\mathbf{P}\bigl[\text{Last}_{[:,0]},\,\text{Last}_{[:,1]},\,N_{d}\bigr]\leftarrow 1$
14:  return $\mathbf{P}$
G.7 Model identifier for directed acyclic graphs
We encode a permutation matrix $\mathbf{P}\in\{0,1\}^{N_{d}\times N_{d}}$ using a compressed Lehmer code consisting of $N_{d}-1$ categorical variables $\{\rho^{\operatorname{cat}}_{1},\ldots,\rho^{\operatorname{cat}}_{N_{d}-1}\}$. Here $\rho^{\operatorname{cat}}_{i}$ has $N_{d}-i+1$ outcomes.
@@ G.7 Model identifier for directed acyclic graphs
Given $\mathbf{P}$ we form an upper-triangular mask $\mathbf{U}\in\{0,1\}^{N_{d}\times N_{d}}$ with zero diagonal. Each entry above the diagonal ($i<j$) is a Bernoulli variable, so $\mathbf{U}$ flattens to $\tfrac{N_{d}(N_{d}-1)}{2}$ bits. The adjacency matrix is $\mathbf{A}=\mathbf{P}^{\top}\mathbf{U}\mathbf{P}$, giving a DAG.
@@
The identifier $\{\rho^{\operatorname{cat}}_{1},\ldots,\rho^{\operatorname{cat}}_{N_{d}-1},\mathbf{U}_{\mathsf{binary}}\}$ is modelled autoregressively by a single $\text{MADE}^{+}$ network, yielding $\mathbf{A}=\mathbf{P}^{\top}\mathbf{U}\mathbf{P}$ upon sampling.
@@
$p(\mathbf{P},\mathbf{U}\mid\gamma)$ $=\frac{1}{N_{d}!}\;\frac{1}{2^{\frac{N_{d}(N_{d}-1)}{2}}}\;\exp\!\bigl(-\gamma\,\mathrm{nEdges}(\mathbf{U})\bigr),$ (56)
$\mathrm{nEdges}(\mathbf{U})$ $=\sum_{i<j}\mathbf{U}_{ij},$ (57)
$\log p$ $=-\log(N_{d}!)-\frac{N_{d}(N_{d}-1)}{2}\log 2-\gamma\,\mathrm{nEdges}(\mathbf{U}).$ (58)
@@
G.8 Neural probability mass function for model indicators over large spaces: $\text{MADE}^{+}$
To represent a distribution over binary strings, we use the Masked Autoencoder for Density Estimation (MADE) [Germain et al., 2015] implementation found in the Durkan et al. [2020] repository. To represent a more complex discrete distribution such as that required by the $\mathbf{P},\mathbf{U}$ representation of a directed acyclic graph, we apply a simple extension to this architecture to allow us to vary the output dimension multiplier. For presentational clarity we call this extension $\text{MADE}^{+}$. The key change in $\text{MADE}^{+}$is the introduction of a per-dimension output multiplier function r(i)\,r(i)\, that determines how many parameters are emitted for the $i$-th input dimension in the autoregressive factorization.